Complete the ordered pair ( , 4) so that it is a solution of x + 2y = 9



Answer by nerdybill(7384)

Complete the ordered pair ( , 4) so that it is a solution of x + 2y = 9
.
Since an "ordered pair" is (x,y)
then the given pair:
( , 4)
says that y=4
.
Plugging the above into:
x + 2y = 9
x + 2(4) = 9
x + 8 = 9
x = 1
.
So, the completed "ordered pair" is:
(1, 4)